OpenStudy (anonymous):

Another inverse function: f(x) = 3 sqrt x- 2

OpenStudy (anonymous):



OpenStudy (anonymous):

this will sound silly, but I know exactly how to do this if that 3 was a cube, not a real number. The real number of 3 is throwing me off.

OpenStudy (anonymous):

this will sound silly, but I know exactly how to do this if that 3 was a cube, not a real number. The real number of 3 is throwing me off.

OpenStudy (anonymous):

f(x) = 3 sqrt ( x- 2) ->( y/3 )^2 = x -2 => x = y^2/9 + 2 Thus f ^-1(x) = x^2/9 + 2

OpenStudy (campbell_st):

let \[x = 3\sqrt{y-2}\] make y the subject \[(x/3)^2 +2 = y\] or \[x^2/9 + 2 = y\]

OpenStudy (anonymous):

awesome, thank you so much. This makes sense now